OpenStudy (shywaifu):

Expand each logorithm. Simplify if possible. Log xyz This is algebra2 and i need help understanding it.

OpenStudy (jonathan34):

picture?

KjSaif (kjsaif):

yea

OpenStudy (jonathan34):

? picture ? please

zepdrix (zepdrix):

Hey you, \(\large\rm \color{salmon}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

We have a log property, which tells us that logs have a way of switching from multiplication to addition,\[\large\rm \log(ab)=\log(a)+\log(b)\]

zepdrix (zepdrix):

It works the same way with 3 terms,\[\large\rm \log(abc)=\log(a)+\log(bc)\]\[\large\rm \log(abc)=\log(a)+\log(b)+\log(c)\]

OpenStudy (shywaifu):

so itd be log (x) + log (y) + log (x) = log xyz

zepdrix (zepdrix):

Yes, that would be the expansion, on the left side of your equals sign. Yay good job!

OpenStudy (shywaifu):

thank you so much!